Donovan & Parents produces soccer shorts and jerseys for youth leagues. Most of the production is done by machine. Data on operations and costs for March follow:

Jerseys Shorts Total
Units produced 10,100 5,100 15,200
Machine-hours used 1,000 830 1,830
Direct labor-hours 250 130 380
Direct materials costs $ 12,100 $ 7,200 $ 19,300
Direct labor costs $ 5,000 $ 2,600 $ 7,600
Manufacturing overhead costs $ 24,476


Management asks the firm’s cost accountant to compute product costs. The accountant first assigns overhead costs to two pools: overhead related to direct materials and overhead related to machine-hours. The analysis of overhead accounts by the cost accountant follows:

Account Amount Related to:
Utilities $ 4,030 Machine-hours
Supplies 2,890 Materials
Machine depreciation and maintenance 10,900 Machine-hours
Purchasing and storing materials 3,286 Materials
Miscellaneous 3,370 Machine-hours


Required:

a. Compute the predetermined overhead rates assuming that Donovan uses machine-hours to allocate machine-related overhead costs and materials costs to allocate materials-related overhead costs. (Round "Machine-related overhead" rate to 2 decimal places and "Materials related overhead" rate as a whole percentage (i.e., 32).)

b. Compute the total costs of production and the cost per unit for each of the two products for March. (Do not round intermediate calculations. Round your "Cost per unit" to 2 decimal places.)
Answer #1

a) Predetermine overhead rate related to machine hour = (4030+10900+3370)/1830 = 10 per machine hour

Predetermine overhead rate related to material cost = 6176/19300 = 32% of material cost

b) calculate total Cost of production and cost per unit

Jerseys Shorts
Direct material 12100 7200
Direct labor 5000 2600
Overhead on machine hour 10000 8300
Overhead on material used 3872 2304
Total Cost of production 30972 20404
Unit 10100 5100
Cost per unit 3.07 4.00